OpenStudy (anonymous):

n ΔABC, BC = a = 16, AC = b = 10, and m∠C = 22°. Which equation can you use to find the value of c = AB?

OpenStudy (baseballguy101):

can i see the answer choices?

OpenStudy (anonymous):

there are none, you have to type your answer

OpenStudy (anonymous):

here is the picture! http://cdn.ple.platoweb.com/PCAP-ASSETS-PROD/f0e6b9bb6f804e60999f40d25bb55b7c

OpenStudy (michele_laino):

here you can apply the theorem of Carnot, namely: \[\Large \begin{gathered} {c^2} = {a^2} + {b^2} - 2ab\cos \left( {22} \right) = \hfill \\ \hfill \\ = {16^2} + {10^2} - 2 \times 16 \times 10\cos \left( {22} \right) = ... \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

i dont get it?

OpenStudy (michele_laino):

ok! Now you have to compute this: \[\Large \sqrt {59.30} = ...?\]

OpenStudy (michele_laino):

what do you get?

OpenStudy (anonymous):

none of it, i've been working on this for 3 days (not this particular question) and it only gets more and more confusing

OpenStudy (michele_laino):

we have: \[\large \begin{gathered} {c^2} = {a^2} + {b^2} - 2ab\cos \left( {22} \right) = \hfill \\ \hfill \\ = {16^2} + {10^2} - 2 \times 16 \times 10\cos \left( {22} \right) = \sqrt {59.30} \cong 7.7 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

am I right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

to see it written out, makes it somewhat easier but when its a different problem and i try and do it myself, i get it wrong

OpenStudy (michele_laino):

that is a simple application of the Theorem of Carnot

OpenStudy (anonymous):

thank you

OpenStudy (michele_laino):

:)